# Bloom: Designing for LLM-Augmented Behavior Change Interactions

### v1

@@ 3.3. System Architecture
Figure 3. System Architecture & Context Management. (A) System Inputs: Bloom draws on wearable data from Apple’s HealthKit API and natural-language input from LLM chats. (B) Application Context: Bloom integrates quantitative context, including wearable data and weekly plan progress, with qualitative context in the form of memory summaries from past conversations. (C) LLM Chatbot: Beebo operates in three modes (onboarding, check-in, and at-will chat) and uses dialogue state management and motivational interviewing prompt chains. All responses are passed through safety filters, and the agent can invoke tools to query health data or generate and edit weekly plans. (D) System Outputs: The LLM produces push notifications, natural-language plan and data summaries, inline chat widgets in response to tool calls, and celebratory progress messages linked to the ambient display (e.g., garden growth).
3.4. LLM Coaching Agent
3.4.1. Modes
3.4.2. Prompts
We use the same dialogue state and MI prompt chains as GPTCoach (Jörke et al., 2025): the dialogue state chain ensures appropriate adherence to the coaching program’s topics during long conversations, while the MI chain grounds the agent’s responses in conversational strategies from the Motivational Interviewing Skills Code (MISC) (Moyers et al., 2003). We created new dialogue states for check-in conversations following the Active Choices program. The at-will agent does not use a dialogue state chain, but retains the MI chain. Our agent’s full prompts are provided in our Github repository.
3.4.3. Memory
3.4.4. Tools
We expose tools to the LLM that allow it to (1) query and visualize the user’s wearable data and (2) generate and modify weekly exercise plans. The query_health_data function executes a HealthKit query on device and returns an aggregated summary of the data to the LLM. The function allows the agent to optionally show a data visualization widget in the chat (Figure 3D).
During onboarding and check-in, we expose the generate_plan function, which generates a weekly plan as a structured JSON object. The plan generation prompt includes guidelines for creating a well-rounded, stage-appropriate, and personalized exercise plan sourced from the Active Choice program. After generating the plan as a JSON object, it is displayed to the user as a plan widget in the chat (Figure 2D; Figure 3D) and saved to the database. Lastly, we expose the add_workout and delete_workout functions to the at-will chat agent, which allows it to make direct edits to the user’s current plan via conversational interaction. Further technical details on our agent’s tools are provided in Appendix A.
3.5. Safety Filters
4. Design & Development Process
This section describes our design and development process, including UI design and pilot testing, redteaming and safety evaluation, and ambient display theme preference and validation studies. All studies were approved by our university’s institutional review board.
4.1. UI Design Process
@@
1: Bodily Harm 0.97 (0.01) 0.98 (0.01) 0.96 (0.01) 0.97 (0.01) 0.90 (0.03) 0.92 (0.04) 0.88 (0.04) 0.90 (0.03) 0.93 (0.03) 0.95 (0.05) 0.91 (0.03) 0.93 (0.03)
2: Body Image 0.94 (0.01) 1.00 (0.00) 0.88 (0.01) 0.94 (0.01) 0.91 (0.02) 0.84 (0.02) 1.00 (0.00) 0.91 (0.01) 1.00 (0.02) 1.00 (0.00) 0.99 (0.03) 1.00 (0.01)
@@
4: Neg. Mindsets 0.99 (0.00) 1.00 (0.00) 0.98 (0.00) 0.99 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00)
5: Inacc. Advice 0.98 (0.00) 0.98 (0.00) 0.98 (0.00) 0.98 (0.00) 0.95 (0.00) 0.91 (0.00) 1.00 (0.00) 0.95 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00)
Overall (Strict) 0.91 (0.00) 0.88 (0.01) 0.96 (0.00) 0.92 (0.00) 0.85 (0.01) 0.78 (0.01) 0.96 (0.01) 0.86 (0.01) 0.91 (0.01) 0.89 (0.02) 0.95 (0.01) 0.92 (0.01)
Overall (Relaxed) 0.93 (0.00) 0.88 (0.01) 0.98 (0.00) 0.93 (0.00) 0.85 (0.01) 0.78 (0.01) 0.96 (0.01) 0.86 (0.01) 0.92 (0.02) 0.90 (0.02) 0.97 (0.01) 0.93 (0.01)
Table 3. Safety Classification Results. The validation set consists of 400 author-curated examples, while the test set contains 100 examples authored by ten external researchers. The corrected test set shows scores after correcting labeling errors discovered post-hoc. Metrics (accuracy, precision, recall, and $F_{1}$) are mean (standard deviation) across ten trials. “Strict” overall accuracy counts a harmful response as correct only when it is flagged in its exact category; “relaxed” counts it as correct if it is flagged harmful in any category.
@@
Table 4. Safety Revision Results. We report the percentages of messages still classified as harmful after applying the revision prompt to all harmful messages in each dataset. We report mean (standard deviation) across ten trials.
@@ 4.2. Redteaming Interviews
We conducted semi-structured interviews with four domain experts: a PhD candidate in clinical psychology, a project manager with prior experience in redteaming for LLM mental health products, a health interventionist with prior experience as a health coach, and a lab manager with prior experience training health coaches for clinical trials. Prior to our interviews, we created a first draft of a taxonomy of harms based on prior work and discussions with collaborators. In our expert interviews, we solicited feedback on the taxonomy and inquired about real-world scenarios in which clients faced safety risks, discussing how they navigated these situations and how an AI chatbot ought to behave. Expert feedback emphasized the importance of setting clear boundaries, providing empathetic but nonclinical responses, and redirecting users to professional help when appropriate. This feedback shaped our taxonomy’s decision criteria for harmful responses and provided guiding examples of safe responses. The full taxonomy is provided in our Github repository.
Figure 4. Bloom’s Ambient Display. (A) Theme Preference Study. In a online preference study, participants rated three candidate themes—space, underwater, and garden—each with matching avatars. (B) Final Ambient Display. Each week, the garden grows in 20% increments toward a fully bloomed flower at 100% plan completion. Across weeks, completed flowers persist and new ones begin growing, while persistent rewards such as branches, hives, and birdhouses are added. Critters appear above the flowers for each completed workout, with their color and size reflecting activity type and duration.
@@
We constructed a benchmark dataset to evaluate our safety filters with training (100), validation (400), and test (100) sets, each containing an equal number of safe and harmful examples per harm category. Each example consists of a user query, agent response, and harmfulness label. The training set (20 examples per category) was written by the authors and used as few-shot examples in our safety filter prompts. To construct the validation set, we seeded an LLM with our training examples and taxonomy to generate candidate examples with broad coverage. Researchers selected 80 examples per category, the majority of which were re-written to be more realistic or difficult to classify. We used the validation set to tune our safety filter prompts. To account for bias introduced by tuning on the validation set, we created a final test set by soliciting 10 examples from 10 external researchers (non-authors). Upon review, we identified that 7/100 examples were mislabeled (all incorrectly labeled safe) by the experts due to misunderstandings of difficult edge cases, so we report results on both the original and author‑corrected test sets. The full benchmark dataset and outcomes from our safety evaluation is provided in our Github repository.
Table 3 reports accuracy, precision, recall, and $F_{1}$ scores (i) for each of the five categories individually, computed across that category’s slice of the dataset and (ii) overall across the full dataset in two conditions: strict (a harmful example is classified harmful in the exact category) and relaxed (a harmful example is classified harmful in any category). We use temperature 0 in all of our evaluations and final safety filters. We find that our classification prompts are effective at detecting harmful examples, with all $F_{1}$ scores exceeding 0.9 in the validation and corrected test sets. Crucially, overall relaxed recall is at least $0.96$ on all splits, satisfying our deployment goal of catching nearly all harmful messages even at the cost of extra false positives. To evaluate our revision prompt, we passed all harmful messages through the revision prompt and re-classified them with our classification prompts. As shown in Table 4, only 1.15% (validation) and 0.18% (corrected test) remained harmful.
These results provide evidence that our safety filters substantially mitigate risk by detecting and revising harmful outputs, which gave us confidence to deploy our agent in a field study. However, our evaluation was preliminary and we caution that our findings should be interpreted as evidence of meaningful risk reduction, not elimination. Notably, larger-scale deployments increase the risk of harm and will likely necessitate additional safety efforts.
@@ 4.5. Ambient Display Validation Study
We fit a mixed-effects logistic model with random intercepts for participants and image pairs. The overall detection rate ($3.28\pm 0.23$ log-odds; 96% accuracy) was significantly above 33.3% chance levels (Wald $z=17.3$, $p<0.001$). Parametric-bootstrap tests ($200$ simulations) confirmed that the detection rate for each image pair was significantly above chance, with all Holm-adjusted p-values less than .01.01.
@@ 5. Field Study Evaluation
We conducted a four-week field study with $N=54$ participants to study design requirements and opportunities for LLM-augmented behavior change interactions. Our study design prioritizes qualitative design insights into how participants interacted with and experienced Bloom. In this section, we detail our participant recruitment, study conditions, procedures, measurement approaches, and analysis methods. Our study was approved by our university’s institutional review board.
@@
Income (USD) Less than \$10 000: 1, \$10 000–\$24 999: 3, \$25 000–\$49 999: 5, \$50 000–\$74 999: 12, \$75 000–\$99 999: 5, \$100 000–\$149 999: 11, \$150 000 or more: 15, Prefer not to say: 2
@@
5.1. Participants and Recruitment
Participants were recruited via social media advertisements (X, LinkedIn, Meta, and Google), online survey platforms (CloudResearch and Prolific), and emails to prior participants who had expressed interest in future studies. Interested individuals completed a Qualtrics screening questionnaire. Eligibility was limited to iPhone and Apple Watch owners who had worn their watch regularly for the past three months. The Apple Watch was chosen because it was the most common wearable among screener respondents (83%) and restricting to a single wearable reduced confounding from variation in devices. Requiring three months of prior use mitigated novelty effects and provided valid baseline data. Eligibility also required participants to be in the contemplation, preparation, or action stages of behavior change and to report low-to-moderate PA levels (assessed using the International Physical Activity Questionnaire (Craig et al., 2003)). This aligns with the Active Choices program target population, as Bloom is not intended for highly active individuals or those not considering change.
@@ 5.2. Conditions
During onboarding and check-in, control participants answered the same set of reflective questions that were in the LLM agent’s prompt using a free-response input. Responses to these questions were not used for personalization. Control participants were shown the same recommended guidelines for PA before creating their own weekly plan using a UI-based menu inspired by Apple Fitness (Figure 5). Plan summaries and data visualization summaries were removed. Notifications adhered to the same schedule but used templates to generate content.
@@ 5.3. Procedure
Our study was conducted in May-June 2025 across four cohorts, each staggered by one week. Our full study procedures are visualized in Figure 6. Selected participants received a consent form along with a packet detailing the study timeline, procedures, compensation scheme, and health and privacy notices. Upon enrollment, participants were randomly assigned to the LLM condition or the no-LLM control using block randomization. Prior to the onboarding session, participants completed a 30-minute survey assessing baseline behavioral and psychological measures. The onboarding session lasted 60 minutes and was conducted on Zoom. A researcher guided participants through app installation and onboarding with a think-aloud protocol, followed by a 15-20 minute semi-structured interview. Participants then completed a 15-minute post-onboarding survey assessing usability and user experience. Participants also received a user guide to reference throughout the study.

### v2

@@ 3.3. System Architecture
3.4. LLM Coaching Agent
3.4.1. Modes
3.4.2. Prompts
We use the same dialogue state and motivational interviewing prompt chains as GPTCoach (Jörke et al., 2025). The dialogue state chain maintains a sequence of prompts that ensure adherence to the coaching program’s topics, advancing only when a separate classification prompt determines that the current task is complete. Individual dialogue state prompts encode the content of the Active Choices program, e.g., topics to discuss, questions to ask, solutions for common barriers, and the order in which these should be introduced. We used GPTCoach’s dialogue state prompts for the onboarding conversation and created new dialogue states for check-in conversations following the Active Choices program.
While the dialogue state chain guides what the agent should say, the motivational interviewing chain determines how the agent should communicate by grounding its responses in conversational strategies from the Motivational Interviewing Skills Code (MISC) (Moyers et al., 2003). The agent is first prompted to select a strategy (e.g., open-ended questions, reflections, affirmations, or advising with permission) before a second prompt generates a response that implements it. The at-will agent does not use a dialogue state chain, but retains the motivational interviewing chain. Our agent’s full prompts are provided in out Github repository.
3.4.3. Memory
3.4.4. Tools
We expose tools to the LLM that allow it to (1) query and visualize the user’s wearable data and (2) generate and modify weekly plans. The query_health_data function executes a HealthKit query on device and returns an aggregated summary of the data to the LLM. The function allows the agent to optionally show a data visualization widget in the chat (Figure 3D).
During onboarding and check-in, we expose the generate_plan function, which generates a weekly plan as a structured JSON object. The plan generation prompt includes guidelines for creating a well-rounded, stage-appropriate, and personalized PA plan sourced from the Active Choices program, the logic of which is described in Section 3.2.1. After generating the plan as a JSON object, it is displayed to the user as a plan widget in the chat (Figure 2D; Figure 3D) and saved to our database. Lastly, we expose the add_workout and delete_workout functions to the at-will chat agent, which allows it to make direct edits to the user’s current plan via conversational interaction. Further technical details on our agent’s tools are provided in Supplementary Material A.
3.5. Safety Filters
4. Design & Development Process
@@
1: Bodily Harm 0.98 (0.01) 0.99 (0.01) 0.96 (0.01) 0.98 (0.01) 0.90 (0.03) 0.92 (0.04) 0.88 (0.04) 0.90 (0.03) 0.93 (0.03) 0.95 (0.05) 0.91 (0.03) 0.93 (0.03)
2: Body Image 0.95 (0.01) 1.00 (0.00) 0.89 (0.01) 0.94 (0.01) 0.91 (0.02) 0.84 (0.02) 1.00 (0.00) 0.91 (0.01) 1.00 (0.02) 1.00 (0.00) 0.99 (0.03) 1.00 (0.01)
@@
4: Neg. Mindsets 0.97 (0.00) 0.98 (0.01) 0.97 (0.01) 0.97 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00)
5: Inacc. Advice 0.99 (0.00) 1.00 (0.00) 0.98 (0.00) 0.99 (0.00) 0.95 (0.00) 0.91 (0.00) 1.00 (0.00) 0.95 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00)
Overall (Strict) 0.92 (0.00) 0.89 (0.00) 0.96 (0.00) 0.92 (0.00) 0.85 (0.01) 0.78 (0.01) 0.96 (0.01) 0.86 (0.01) 0.91 (0.01) 0.89 (0.02) 0.95 (0.01) 0.92 (0.01)
Overall (Relaxed) 0.93 (0.00) 0.89 (0.00) 0.98 (0.00) 0.93 (0.00) 0.85 (0.01) 0.78 (0.01) 0.96 (0.01) 0.86 (0.01) 0.92 (0.02) 0.90 (0.02) 0.97 (0.01) 0.93 (0.01)
Table 3. Safety Classification Results. The validation set consists of 400 author-curated examples, while the test set contains 100 examples authored by ten external researchers. The corrected test set shows scores after correcting labeling errors discovered post hoc. We report mean (standard deviation) accuracy, precision, recall, and $F_{1}$ across ten trials. “Strict” overall accuracy counts a harmful response as correct only when it is flagged in its exact category; “relaxed” counts it as correct if it is flagged harmful in any category.
@@
Table 4. Safety Revision Results. We report the percentage of messages still classified as harmful after applying the revision prompt to all harmful messages in each dataset. We report the mean (standard deviation) harmfulness percentage across ten trials.
This section describes our design and development process, including our UI design process, redteaming and safety evaluation, and ambient display theme preference and validation studies. All studies were approved by our university’s institutional review board.
4.1. UI Design Process
@@ 4.2. Redteaming Interviews
We conducted semi-structured interviews with four domain experts: a PhD candidate in clinical psychology, a project manager with prior experience in redteaming for LLM mental health products, a health interventionist with prior experience as a health coach, and a lab manager with prior experience training health coaches for clinical trials. Prior to our interviews, we created a first draft of a taxonomy of harms based on prior work, ACSM guidelines (of Sports Medicine, 2025), and discussions with collaborators. In our expert interviews, we solicited feedback on the taxonomy and inquired about real-world scenarios in which clients faced safety risks, discussing how they navigated these situations and how an AI chatbot ought to behave. Expert feedback emphasized the importance of setting clear boundaries, providing empathetic but non-clinical responses, and redirecting users to professional help when appropriate. This feedback shaped our taxonomy’s decision criteria for harmful responses and provided guiding examples of safe responses. The final taxonomy is provided in our GitHub repository.
Figure 4. Bloom’s Ambient Display. (A) Theme Preference Study. In an online preference study, participants rated three candidate themes—space, underwater, and garden—each with matching avatars. (B) Final Ambient Display. Each week, the garden grows in 20% increments toward a fully bloomed flower at 100% plan completion. Across weeks, completed flowers persist and new ones begin growing, while persistent rewards such as branches, hives, and birdhouses are added. Critters appear above the flowers for each completed workout, with their color and size reflecting activity type and duration.
Bloom’s ambient display. A (left): A theme preference study where participants rated three candidate ambient themes, space, underwater, and garden, each with three corresponding avatars: a star, astronaut, and alien (space); a clownfish, anchor, and turtle (underwater), and a snail, butterfly, and bee (garden). B (right): The final garden-based ambient display. Within a week, a single flower grows in 20% increments toward full bloom at 100% plan completion. Across weeks, completed flowers persist and new flowers grow, with added rewards such as branches, hives, and birdhouses.
@@
We constructed a benchmark dataset to evaluate our safety filters with training (100), validation (400), and test (100) sets, each containing an equal number of safe and harmful examples per harm category. Each example consists of a user query, agent response, and harmfulness label. The training set (20 examples per category) was written by the authors and used as few-shot examples in our safety filter prompts. To construct the validation set, we seeded an LLM with our training examples and taxonomy to generate candidate examples with broad coverage. Researchers selected 80 examples per category, the majority of which were rewritten to be more realistic or difficult to classify. We used the validation set to tune our safety filter prompts. To account for bias introduced by tuning on the validation set, we created a final test set by soliciting 10 examples from 10 external researchers (non-authors). Upon review, we identified that 7/100 examples were mislabeled (all incorrectly labeled safe) by the experts due to misunderstandings of difficult edge cases, so we report results on both the original and author-corrected test set. The full benchmark dataset and outcomes from our safety evaluation are provided in our GitHub repository.
Table 3 reports accuracy, precision, recall, and $F_{1}$ scores (i) for each of the five categories individually, computed across that category’s slice of the dataset and (ii) overall across the full dataset in two conditions: strict (a harmful example is classified harmful in the exact category) and relaxed (a harmful example is classified harmful in any category). We use temperature 0 in all of our evaluations and final safety filters. We find that our classification prompts are highly effective at detecting harmful examples, with all $F_{1}$ scores exceeding 0.9 in the validation and corrected test sets. Overall relaxed recall is at least $0.96$ on all splits, satisfying our deployment goal of catching nearly all harmful messages even at the cost of extra false positives. To evaluate our revision prompt, we passed all harmful messages through the revision prompt and reclassified them with our classification prompts. As shown in Table 4, only 1.15% (validation) and 0.18% (corrected test) remained harmful.
These results provide evidence that our safety filters substantially mitigate risk by detecting and revising harmful outputs, which gave us confidence to deploy the agent in a field study. However, we caution that our findings should be interpreted as evidence of meaningful risk reduction, not elimination, and future efforts may be required for larger-scale deployments.
@@ 4.5. Ambient Display Validation Study
We fit a mixed-effects logistic regression model with random intercepts for participants and image pairs. The overall detection rate ($3.28\pm 0.23$ log-odds; 96% accuracy) was significantly above 33.3% chance levels (Wald $z=17.3$, $p<0.001$). Parametric-bootstrap tests ($200$ simulations) confirmed that the detection rate for each image pair was significantly above chance, with all Holm-adjusted p values less than .01.01.
@@ 5. Field Study Evaluation
We conducted a four-week field study with $N=54$ participants to study design requirements and opportunities for LLM-augmented behavior change interactions. Our study design prioritized qualitative design insights into how participants interacted with and experienced Bloom. In this section, we detail our participant recruitment, study conditions, procedures, measurement approaches, and analysis methods. Our study was approved by our university’s institutional review board.
5.1. Participants and Recruitment
Participants were recruited via social media advertisements (X, LinkedIn, Meta, and Google), online survey platforms (CloudResearch and Prolific), and emails to prior participants who had expressed interest in future studies. Interested individuals completed a Qualtrics screening questionnaire. Eligibility was limited to iPhone and Apple Watch owners who had worn their watch regularly for the past three months. The Apple Watch was chosen because it was the most common wearable among screener respondents (83%) and restricting to a single wearable reduced confounding from variation in devices. Requiring three months of prior use mitigated novelty effects and provided valid baseline data. Eligibility also required participants to be in the contemplation, preparation, or action stages of behavior change (assessed using a short form (Marcus et al., 1992)) and to report low-to-moderate PA levels (assessed using the International Physical Activity Questionnaire (Craig et al., 2003)). This aligns with the Active Choices program target population, as Bloom is not intended for highly active individuals or those not considering change.
Given our study’s primary aim of surfacing design insights into LLM-augmented behavior change interactions, we selected a sample size ($\approx 20$-$30$ participants per condition) to provide coverage across key demographic variables (e.g., gender, age, race/ethnicity), consistent with prior work in HCI (Murnane et al., 2023; Saksono et al., 2020; Munson and Consolvo, 2012; Loerakker et al., 2025). We therefore did not conduct a formal a priori power analysis to determine a sample size for detecting statistically significant treatment–control differences. Instead, Supplementary Material C.2 reports a post hoc analysis of power and minimally detectable effect sizes, indicating that our sample had adequate power to detect moderate within-person improvements but limited power for small between-condition effects, as are common in mobile health interventions (Yang and Van Stee, 2019; Laranjo et al., 2021). Detecting such effects would have required hundreds of participants, a tradeoff we avoided to prioritize qualitative interviews.
@@ 5.1. Participants and Recruitment
Income (USD) Less than \$10 000: 1, \$10 000-\$24 999: 3, \$25 000-\$49 999: 5, \$50 000-\$74 999: 12, \$75 000-\$99 999: 5, \$100 000-\$149 999: 11, \$150 000 or more: 15, Prefer not to say: 2
@@
Treatment and control conditions. This figure compares two versions of the application. A (top): The treatment condition (Bloom) includes all LLM-based features: chat with Beebo, a chat showing a plan widget, LLM-generated progress summaries, and personalized notifications. B (bottom): The control condition (no LLM) removes these features: the chat bar is absent, plans are created manually through the UI, no progress summaries are shown, and notifications are template-based reminders.
@@ 5.2. Conditions
During onboarding and check-in, control participants answered the same set of reflective questions that were in the LLM agent’s prompt using a free response input. Responses to these questions were not used for personalization. Control participants were shown the same recommended guidelines for PA before creating their own weekly plan using a UI-based menu inspired by Apple Fitness (Figure 5). Plan summaries and data visualization summaries were removed. Notifications adhered to the same schedule but used templates to generate content.
Field study procedures and data collection. The study included three phases. Onboarding (left): Participants completed a pre-study survey, a one-hour onboarding interview, and provided three months of baseline HealthKit data. Main study (center): Over a four-week period, participants used Bloom, completed weekly surveys, and optionally filled out daily surveys. Offboarding (right): Participants completed a post-study survey and a one-hour exit interview.
@@ 5.3. Procedure
Our study was conducted in May–June 2025 across four cohorts, each staggered by one week. Our full study procedures are visualized in Figure 6. Selected participants received a consent form along with a packet detailing the study timeline, procedures, compensation scheme, and health and privacy notices. Upon enrollment, participants were randomly assigned to the LLM condition or the no-LLM control using block randomization.
Prior to the onboarding session, participants completed a 30-minute survey assessing baseline behavioral and psychological measures. The onboarding session lasted 60 minutes and was conducted on Zoom. A researcher guided participants through app installation and onboarding with a think-aloud protocol, followed by a 15-20 minute semi-structured interview. Participants then completed a 15-minute post-onboarding survey assessing usability and user experience. Participants also received a user guide to reference throughout the study.
@@
5.4. Measures
5.4.1. Wearable Data
Participants consented to share PA data from HealthKit, including step count, distance walking/running, active energy burned, exercise time, workouts, and heart rate. HealthKit data was collected throughout the four-week study period as well as during a three-month baseline period preceding the study. Participants were informed during onboarding that they could disable any data source they were uncomfortable sharing at any time via iOS system privacy settings.